### v1

@@ Abstract
We present LidarDM, a novel LiDAR generative model capable of producing realistic, layout-aware, physically plausible, and temporally coherent LiDAR videos. LidarDM stands out with two unprecedented capabilities in LiDAR generative modeling: (i) LiDAR generation guided by driving scenarios, offering significant potential for autonomous driving simulations, and (ii) 4D LiDAR point cloud generation, enabling the creation of realistic and temporally coherent sequences. At the heart of our model is a novel integrated 4D world generation framework. Specifically, we employ latent diffusion models to generate the 3D scene, combine it with dynamic actors to form the underlying 4D world, and subsequently produce realistic sensory observations within this virtual environment. Our experiments indicate that our approach outperforms competing algorithms in realism, temporal coherency, and layout consistency. We additionally show that LidarDM can be used as a generative world model simulator for training and testing perception models.
Abstract
In the following supplementary material, we provide additional model details and ablations (Sec. 1). In Sec. 2 we provide additional qualitative LidarDM results, including a baseline comparison, a close look at pedestrian details, and visualizations of map condition alignment. In Sec. 3, we provide additional downstream applications such as crafting safety-critical and out-of-distribution scenarios for self driving. In Sec. 3.3, we perform a Sim2Real data augmentation experiment on an end-to-end LiDAR-based planner.
Keywords:
LiDAR Generation Scene Generation Self-driving
1 Introduction
Figure 1: We present LidarDM, a novel 4D LiDAR generative model. Our generated LiDAR videos simultaneously enjoy the benefits of being realistic, layout-conditioning, physically plausible, diverse, and temporally coherent.
Generative models have become notable in understanding data distributions and content creation, e.g. in image and video generation [33, 54, 53, 55, 52, 10], 3D object generation [19, 52, 10, 38], compression [5, 68, 29], and editing [47, 37]. Generative models also show significant promise for simulation [82, 46, 11, 66, 18, 34, 64, 60, 6, 76], capable of creating realistic scenarios and their associated sensory data for training and evaluating safety-critical embodied agents, such as in robotics and autonomous vehicles, without the need of expensive manual modeling of the real world. These capabilities are crucial for applications that rely on extensive closed-loop training or scenario testing.
@@ 1 Introduction
While advancements in conditional image and video generation [35, 27, 44, 15] have been remarkable, the specific task of generatively creating scenario-specific, realistic LiDAR point cloud sequences for autonomous driving application remains under-explored. Current LiDAR generation methods fall into two broad categories, each of which suffers from specific challenges:
  1. 1.
    LiDAR generative modeling methods [83, 72, 8, 79] are currently limited to single-frame generation and do not provide the means for semantic controllability and temporal consistency.
  2. 2.
    LiDAR resimulation [14, 65, 17, 46, 74, 67] depends heavily on user-created or real-world collected assets. This induces a high operational cost, restricts diversity, and limits broader applicability.
@@
To address these challenges, we propose LidarDM (Lidar Diffusion Model), which creates realistic, layout-aware, physically plausible, and temporally coherent LiDAR videos. We explore two novel capabilities that have not been previously addressed: (i) LiDAR synthesis guided by driving scenarios, which holds immense potential for simulation in autonomous driving, and (ii) 4D LiDAR point cloud synthesis aimed at producing realistic and temporally coherent sequences of labeled LiDAR point clouds. Our key insight to achieving these goals lies in first generating and composing the underlying 4D world and then creating realistic sensory observations within this virtual environment. To achieve this, we integrate existing 3D object generation approaches to create dynamic actors and develop a novel approach for large-scale 3D scene generation based on the latent diffusion model. This method produces realistic and diverse 3D driving scenes from a coarse semantic layout, which to our knowledge, is one of the first of its kind. We apply trajectory generation to create dynamic effects while ensuring authentic interactions among actors and between actors and the scene. Finally, we compose the 3D world at each time step and perform stochastic raycasting simulation to produce the final 4D LiDAR sequence. As shown in Fig. 1, our generated results are diverse, align with the layout conditions, and are both realistic and temporally coherent.
@@
Our experimental results demonstrate that individual frames generated by LidarDM exhibit realism and diversity, with performance that is on-par with state-of-the-art techniques in unconditional single-frame LiDAR point cloud generation. Moreover, we show that LidarDM can produce LiDAR videos that maintain temporal coherency, outperforming a robust stable diffusion sensor generation baseline. To our knowledge, this is the first LiDAR generative method with this capability. We further demonstrate LidarDM’s conditional generation by showing that the generated LiDAR matches well with ground-truth LiDAR on matching map conditions. Lastly, we illustrate that the data generated by LidarDM exhibit a minimal domain gap when tested with perception modules trained on real data and can also be used to augment training data to significantly boost the performance of 3D detectors. This gives premise for using generative LiDAR models to create realistic and controllable simulations for training and testing driving models (Please refer to Sec. 2 and Fig. 2 for detailed applications of LidarDM).
2 Related Works
Figure 2: Applications of LidarDM: (a) generating LiDAR that aligns well with the map (color boxes highlight the consistency between the lidar and the map) without 3D capturing or modeling; (b) providing sensor data for an existing traffic simulator (Waymax [20]), enabling safety-critical scenarios evaluation from pure sensor data; (c) generate large volume Lidar data with controllable obstacles locations (treated as ground-truth labels, which are free to obtain) to improve perception models via pre-training without expensive data capturing and labelling.
LiDAR Simulation.
@@ 2 Related Works
Realistic LiDAR sensor simulation is crucial for robotics and self-driving vehicle training and testing. Traditional LiDAR generation methods use raycasting-based physical approaches. Simulators like CARLA [14] and AirSim [59] create environments with static (buildings, trees, street lights) and dynamic objects (cars, bicycles, buses). In these settings, virtual LiDAR sensors are placed, casting rays to calculate depth through ray-triangle intersections. Such approaches are simple and easy to integrate, hence are widely used in robot simulation [45, 4]. Asset-based physical simulation methods for LiDAR face limitations in realism and scalability due to three key issues: the need for 3D assets, which are costly and limit variations; further, they also face challenges in closing the sim2real gap for both asset design and physics simulation.
Recent research attempts to address these shortcomings through data-driven approaches. LiDARSim [46] uses collected LiDAR sequences to reconstruct maps and dynamic assets. Subsequent works have improved asset reconstruction using neural fields-based approaches like NeRF [30, 67] or neural feature fields [74], and with automatic alignment [58], or with past LiDAR readings [30]. Additionally, realistic physics effects such as ray dropping [49] and snow [21] have also been modeled. However, constructing detailed 3D maps and objects from real-world sensor data is often costly and not scalable, typically requiring multiple passes over the same locations.
Our approach, unlike traditional LiDAR simulations, is purely generative, eliminating the need for man-made or reconstructed assets and allowing for easy creation of numerous virtual worlds. In Fig. 2 (a), we present a realistic generated LiDAR point cloud of Champs-Élysées from only a hand-crafted map layout (without any actual sensor data from France), which no re-simulation methods can achieve.
LiDAR Generation.
Generative models provide a promising alternative for creating realistic LiDAR point clouds without reconstructing real-world environments. Early LiDAR generation work exploited the range image representation for LiDAR generation. The pioneering work by Caccia et al. [8] focused on using GANs and VAEs for unconditional generation and performing reconstruction tasks for noisy LiDAR readings. LiDARGen [83] showed that using a score-based diffusion model provides improved generative capability and can be used in downstream classifier-guided sampling tasks such as point cloud upsampling. The range image representation offers the benefit of physically accurate rendering at the cost of being ego-centric, rather than scene-centric. Recently, UltraLiDAR [72] proposed using a BEV voxel grid representation for LiDAR generation. Sampling is performed with a VQVAE [50] in a learned discrete latent space, ensuring a dense latent space. The BEV-centric representation provides the benefit of improved layout coherence and metrics, at the cost of not guaranteeing a physically accurate LiDAR sample. However, despite significant advancements, current generative models do not support conditioning on semantic layouts. This omission makes the generation process less controllable, reducing its practical applications. Additionally, a model ensuring temporal consistency in LiDAR video generation remains to be developed.
Unlike previous methods, our model addresses this challenge by (1) generalizing the task of LiDAR generation to geometry generation and data-driven ray casting in a novel field representation; (2) guiding generation with a BEV HD Map layout; and (3) allowing complete control over dynamic scene composition by adding generated 3D models and enabling free movement of the virtual LiDAR sensor within the frame. Fig. 2 (b) and (c), respectively, show that LidarDM’s point cloud can provide realistic LiDAR information to an existing traffic simulator for autonomous driving’s safety-critical case evaluation thanks to its temporal consistency, and can improve perception accuracy (Sec. 4.5) thanks to its full controllability of scenarios, which means ground-truth detection labels are free to obtain. These two important benefits cannot be achieved with any other LiDAR generative methods.
Diffusion Models.
Our model builds upon the recent advancement in latent diffusion models. Directly applying diffusion models [24, 62] on data can be burdensome on the denoising network due to the issue of data sparsity. Latent diffusion alleviates this issue by performing the diffusion process on a dense latent space of an autoencoder. Models such as stable diffusion [54], stable diffusion XL [51], and MaskGIT [9] have championed latent diffusion as a highly effective technique for generative modeling.
Sampling: A key challenge in diffusion-based modeling lies in selecting the sampling procedure. Early methods utilized Langevin dynamics or ancestral sampling. Recent sampling quality and speed improvements have come from deterministic non-Markovian techniques like DDIM [61] and PNDM [40]. Additionally, the differential equation interpretation of diffusion models has led to the development of samplers like Euler [32] and DPM-Solver [43]. We empirically find Euler works well in practice for our model.
Conditioning: Various strategies have been devised to integrate conditions into the diffusion model generation process. Classifier-based guidance (or posterior sampling) [62, 47, 83] utilizes the gradient of a classifier to enhance the model’s score function. Classifier-free guidance [23] offers a method for training and sampling class-conditioned diffusion models. Controlnet [80] introduces a technique for adding controllability to a pre-trained diffusion model through a class-conditioned hypernetwork. Our approach leverages classifier-free guidance.
In diffusion-based video generation [25, 31, 71, 70, 22, 73], maintaining consistency is a central challenge. Various approaches have been proposed, leveraging motion modeling, interpolation, or batch sampling to enhance consistency. Our model is the first to focus on LiDAR video generation. Unlike other methods, our technique capitalizes on LiDAR’s unique attribute of underlying 3D world alignment, significantly improving temporal consistency.
Figure 3: Overview of the LidarDM: Given the input traffic layout at time $t=0$, LidarDM begins by generating actors and the static scene. We then generate the motion of the actors and the egocar, and compose the underlying 4D world. Finally, a generative- and physics-based simulation is used to create realistic 4D sensor data.
3 Layout-Guided LiDAR Video Generation
@@
Our goal is to create a realistic, physically plausible, and temporally consistent LiDAR sequence that enables a free viewpoint based on a given bird’s eye view semantic layout in a purely generative manner without relying on any pre-collected assets like 3D maps. To our knowledge, this is the first solution of its kind, addressing layout-conditioned LiDAR generation and LiDAR video generation. The key to achieving this lies in first generating and composing the underlying 3D world, followed by using generative simulation to create realistic sensory observations. We begin by formulating the generation as a joint 4D scene generation task (Sec. 3.1). Next, we discuss leveraging 3D diffusion models to create static and dynamic elements, ensuring their faithful interaction (Sec. 3.2). Finally, a sensor generation procedure is executed to produce the final LiDAR video (Sec. 3.3). Fig. 3 depicts the overview of our method.
3.1 Problem Formulation
@@
Formally, given an input layout $\mathcal{I}\in{\mathbb{R}}^{L\times W\times M}$ representing traffic elements from a bird’s eye view (where $L$, $W$, and $M$ are length, width, and map classes, respectively), our goal is to generate a LiDAR point cloud video $\mathcal{X}=\{\mathbf{x}_{t}\}$, with each $\mathbf{x}_{t}\in{\mathbb{R}}^{N\times 3}$ being a point cloud at frame $t$ with $\mathbf{x}_{0}$ matching the input layout. This conditional generation setting offers full controllability, and hence lays the foundation for a practical asset-free simulator. Note that in the absence of a map, our approach defaults to unconditional generation.
4D World Representation.
@@
Our key technical innovation to address the challenge lies in jointly modeling the generation of underlying 4D world together sensor generation. We define the world scene representation as $\mathcal{W}=\{\mathbf{s},\{\mathbf{o}_{i}\}_{i=0}^{N}\}$, where $\mathbf{s}$ represents a static scene geometry and $\mathbf{o}_{0},...,\mathbf{o}_{N}$ are dynamic objects. Both are represented in the form of an occupancy grid. To model dynamics, we additionally consider the actions of these dynamic objects in the form of trajectories $\mathcal{P}=\{\boldsymbol{\tau}_{0},...,\boldsymbol{\tau}_{T}\}$, with $\boldsymbol{\tau}_{t}=\{\boldsymbol{\xi}_{\mathrm{ego}},\{\boldsymbol{\xi}_{i,t}\}_{i=0}^{N}\}$ representing the pose of actor $i$ at time $t$ as well as egocar pose $\boldsymbol{\xi}_{\mathrm{ego}}$. The pose for rigid objects and the egocar lies in the $\mathbb{SE}(3)$ space, while for articulated objects like pedestrians, it is represented as a kinematic chain. A composed scene represent the states of the world at $t$, incorporating the poses of the ego car and dynamic objects at time $t$, is denoted by $\mathcal{W}_{t}=\pi(\mathcal{W},\boldsymbol{\tau}_{t})$, where $\pi$ is a composition operator applying transformations to each actor.
4D World and LiDAR Generation.
@@
$\underbrace{p(\mathbf{s}|\mathcal{I})\cdot\prod_{i}p(\mathbf{o}_{i}|\mathcal{I})}_{\text{3D scene and object gen}}\cdot\underbrace{\prod_{t}p(\boldsymbol{\tau}_{t}|\boldsymbol{\tau}_{<t},\mathcal{W},\mathcal{I})}_{\text{trajectory gen}}\cdot\underbrace{\prod_{t}p(\mathbf{x}_{t}|\boldsymbol{\tau}_{t},\mathcal{W})}_{\text{sensor simulation}}.$ (1)
@@
Figure 4: Our 3D scene generation pipeline. First, accumulated point clouds are used to reconstruct each ground truth mesh sample. Next, a variational autoencoder is trained to compress meshes into a latent code. Finally, a map-conditioned diffusion model is trained to perform sampling within the latent space of the VAE, yielding novel samples.
3.2 Scene, Object and Trajectory Generation
We decompose the world into a static background scene, constant over time, and dynamic foreground objects that move. This decomposition simplifies the challenging 4D world generation into more manageable tasks: creating object geometries and generating dynamic effects. This modeling approach ensures temporal consistency (e.g. keeping cars’ shapes constant and walls and trees remain still over time) and physical plausibility (e.g. ensuring correct occlusion reasoning).
Scene Generation.
The scene generation addresses the problem of sampling the geometry of a scene from a given input layout $\mathcal{I}$: $\mathbf{s}\sim p(\mathbf{s}|\mathcal{I})$. We parameterize the 3D scene using a signed distance field, $\mathbf{s}\in\mathbb{R}^{L\times W\times H}$, where each entry $s_{j}$ encodes the truncated signed distance to the surface, $-1\leq s_{j}\leq 1$, with negative values indicating outside the mesh and positive values inside the surface.
Model: We leverage the latent diffusion model [54, 54] to tackle this challenge of modeling and sampling from $p(\mathbf{s}|\mathcal{I})$. We choose the latent diffusion model for its capacity to sample high-quality data while effectively incorporating strong conditional guidance. Specifically, our model encodes the high-dimensional SDF volume $\mathbf{s}$ into a continuous latent representation $\mathbf{z}$ using an encoder-decoder structure [54] with a scene encoder $E_{\theta}(\mathbf{s})=\mathbf{z}$ and a scene decoder $D_{\theta}(\mathbf{z})=\tilde{\mathbf{s}}$. This encoder-decoder structure efficiently compresses the input data into a lower-dimensional latent space, enabling more effective and efficient sampling. Additionally, we encode our high-definition map layout $\mathcal{I}$ into a latent space $\mathbf{c}=M_{\theta}(\mathcal{I})$, allowing for more compact conditioning.

### v2

@@ Abstract
We present LidarDM, a novel LiDAR generative model capable of producing realistic, layout-aware, physically plausible, and temporally coherent LiDAR videos. LidarDM stands out with two unprecedented capabilities in LiDAR generative modeling: (i) LiDAR generation guided by driving scenarios, offering significant potential for autonomous driving simulations, and (ii) 4D LiDAR point cloud generation, enabling the creation of realistic and temporally coherent sequences. At the heart of our model is a novel integrated 4D world generation framework. Specifically, we employ latent diffusion models to generate the 3D scene, combine it with dynamic actors to form the underlying 4D world, and subsequently produce realistic sensory observations within this virtual environment. Our experiments indicate that our approach outperforms competing algorithms in realism, temporal coherency, and layout consistency. We additionally show that LidarDM can be used as a generative world model simulator for training and testing perception models. We release our source code and checkpoints at https://github.com/vzyrianov/LidarDM
I INTRODUCTION
Generative models are notable in areas such as image and video generation [1, 2, 3, 4], 3D generation [5, 6, 7], compression [8, 9], and editing [10, 11]. More recently, they have shown great promise in robotics by generating realistic scenarios and sensory data for training and validating embodied agents, significantly reducing the cost of real-world experiments [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@ I INTRODUCTION
While advancements in conditional image and video generation [24, 25, 26, 27] have been remarkable, the specific task of generatively creating scenario-specific, realistic LiDAR point cloud sequences for autonomous driving application remains under-explored. Current LiDAR generation methods fall into two broad categories, each of which suffers from specific challenges. (i) LiDAR generative modeling methods [28, 29, 30, 31] are currently limited to single-frame generation and do not provide the means for semantic controllability and temporal consistency. (ii) LiDAR resimulation [32, 33, 34, 13, 35, 36] relies on user-created or real-world collected assets, this induces a high cost, restricts diversity, and limits broader applicability.
Figure 1: LiDAR sequences generated by LidarDM are realistic, layout-conditioned, physically plausible, and temporally coherent. We show 3 frames of 2 different videos generated from the same layout. In the layout visualization, the red rectangles are dynamic vehicles, gray rectangles are static vehicles, blue dots are pedestrians, the green dot is the ego vehicle sensor, and the gray curves are road boundaries and lane markings.).
@@
To address these challenges, we propose LidarDM (Lidar Diffusion Model), which creates realistic, layout-aware, physically plausible, and temporally coherent LiDAR videos. LidarDM enables two previously unexplored capabilities : (i) controllable LiDAR synthesis guided by driving scenarios, which holds immense potential for simulation in autonomous driving, and (ii) 4D LiDAR synthesis for generating realistic and temporally coherent sequences of labeled LiDAR point clouds. Our key insight lies in first generating and composing the underlying 4D world and then creating realistic sensory observations within this virtual environment. We develop a novel approach for large-scale 3D scene generation based on the latent diffusion model and integrate existing 3D object generators for dynamic actors. This method produces realistic and diverse 3D driving scenes from a coarse semantic layout, which to our knowledge, is one of the first of its kind. Finally, we simulate dynamic and realistic driving scenario, compose the 3D world at each time step, and perform stochastic raycasting to produce the final 4D LiDAR sequence. Our generated results are diverse, realistic, temporally coherent, and align with the layout (Fig. 1).
@@
Our experimental results demonstrate that individual frames generated by LidarDM exhibit realism and diversity, with performance on-par with state-of-the-art techniques in unconditional single-frame LiDAR point cloud generation. Moreover, we show that LidarDM can produce temporally coherent LiDAR videos, outperforming a robust stable diffusion baseline. We further demonstrate LidarDM’s conditional generation by showing that the generated LiDAR matches well with ground-truth LiDAR on matching map conditions for both geometry and intensity. Among generative LiDAR methods, LidarDM is the first to support map-conditioned generation and the first to support sequence generation. Lastly, we illustrate that the data generated by LidarDM exhibit a minimal domain gap when tested with perception modules trained on real data and can also be used to augment training data to significantly improve 3D detectors and planners. This gives premise for using generative LiDAR models to create realistic and controllable simulations for training and testing driving models. For detailed applications of LidarDM, please refer to Sec. II and Fig. 2.
II Related Works
II-1 LiDAR Simulation
@@
Realistic LiDAR sensor simulation is crucial for robotics and self-driving vehicle training and testing. Simulators like CARLA [32] and AirSim [37] create environments with static (buildings, trees, street lights) and dynamic (cars, bicycles, buses) assets and simulate LiDAR with raycasting. Such approaches are simple and easy to integrate, hence are widely used in robot simulation [38, 39]. However, these methods face limitations in realism and scalability due to two key issues: (i) the need for 3D assets, which are costly and limit variations; (ii) the sim2real gap for both asset design and physics simulation.
Recent data-driven approaches reconstruct objects and environments from real-world LiDAR as meshes [13, 40] or neural fields [41, 36] from which LiDAR are simulated from raycasting or volume rendering. However, these methods require rendering LiDAR from models constructed from real-world scenes, which is costly and not scalable. In contrast, LidarDM creates scenes in a purely generative manner, eliminating the need for man-made or reconstructed assets and environments, and can generate point clouds from environments unseen during training. For example, Champs-Élysées (in Fig. 2 (a)) was created from a hand-crafted map layout, which no re-simulation method can achieve.
II-2 LiDAR Generation
Generative models provide a promising alternative for creating realistic LiDAR point clouds without reconstructing real-world environments. Early LiDAR generation works utilized the range image representation using GANs [30], VAEs [30], and diffusion models [28, 42, 43, 44]. Later works used voxel representations with VQGANs [29]. However, these methods only generate single frame LiDAR, and do not provide controllable or video generation. LidarDM addresses these issues by conditionally generating a 4D world and performing physics-based raycasting. As shown in Fig.2 (b) and (c), LidarDM provides realistic LiDAR data to traffic simulators thanks to its temporal consistency and can train perception models (Sec.IV-E) with paired semantic layout and generated LiDAR—benefits unmatched by other generative methods.
III Layout-Guided LiDAR Video Generation
Figure 2: Applications of LidarDM: (a) generating layout-conditioned LiDAR (color boxes highlight the lidar-map consistency); (b) creating sensor data for traffic simulator [45] to enable end-to-end safety-critical scenarios evaluation; (c) generate large volume Lidar data with known ground truth labels to improve perception models without data capturing and labelling.
@@ III Layout-Guided LiDAR Video Generation
Our goal is to create a realistic, physically plausible, and temporally consistent LiDAR sequence that enables a free viewpoint based on a given bird’s eye view semantic layout in a purely generative manner without relying on any pre-collected assets like 3D maps. The key to achieving this lies in first generating and composing the underlying 3D world, followed by using generative simulation to create realistic sensory observations. We begin by formulating generation as a joint 4D scene generation task (Sec. III-A). Next, we discuss leveraging 3D diffusion models to create static and dynamic elements, and ensuring faithful interactions (Sec. III-B). Finally, a sensor generation procedure is executed to produce the final LiDAR video (Sec. III-C). Fig. 3 depicts the overview of our method.
III-A Problem Formulation
@@
Formally, given an input layout $\mathcal{I}\in{\mathbb{R}}^{L\times W\times M}$ representing traffic elements from a bird’s eye view where $L$, $W$, and $M$ are length, width, and map classes (lanes, roads, crosswalks), respectively, our goal is to generate a LiDAR point cloud video $\mathcal{X}=\{\mathbf{x}_{t}\}$, with each $\mathbf{x}_{t}\in{\mathbb{R}}^{N\times 4}$ being a point cloud (3D location and intensity) at frame $t$ with $\mathbf{x}_{0}$ matching the input layout. This conditional generation setting offers full controllability, and hence lays the foundation for a practical asset-free simulator. Without a map, our approach defaults to unconditional generation (i.e., in Sec. IV-C).
III-A1 4D World Representation
@@
Our key technical innovation to address the challenge lies in jointly modeling the generation of underlying 4D world together with sensor generation. We define the world scene representation as $\mathcal{W}=\{\mathbf{s},\{\mathbf{o}_{i}\}_{i=0}^{N}\}$, where $\mathbf{s}$ represents a static scene geometry and $\mathbf{o}_{0},...,\mathbf{o}_{N}$ are dynamic objects. Both are represented in the form of an occupancy grid. To model dynamics, we additionally consider the actions of these dynamic objects in the form of trajectories $\mathcal{P}=\{\boldsymbol{\tau}_{0},...,\boldsymbol{\tau}_{T}\}$, with $\boldsymbol{\tau}_{t}=\{\boldsymbol{\xi}_{\mathrm{ego}},\{\boldsymbol{\xi}_{i,t}\}_{i=0}^{N}\}$ representing the pose of actor $i$ at time $t$ as well as egocar pose $\boldsymbol{\xi}_{\mathrm{ego}}$. The pose for rigid objects and the egocar lies in the $\mathbb{SE}(3)$ space, while for articulated objects like pedestrians, it is represented as a kinematic chain. A composed scene represent the states of the world at $t$, incorporating the poses of the ego car and dynamic objects at time $t$, is denoted by $\mathcal{W}_{t}=\pi(\mathcal{W},\boldsymbol{\tau}_{t})$, where $\pi$ is an operator composing actors to world.
III-A2 4D World and LiDAR Generation
@@
$\underbrace{p(\mathbf{s}|\mathcal{I})\cdot\prod_{i}p(\mathbf{o}_{i}|\mathcal{I})}_{\text{3D scene and object gen}}\cdot\underbrace{\prod_{t}p(\boldsymbol{\tau}_{t}|\boldsymbol{\tau}_{<t},\mathcal{W},\mathcal{I})}_{\text{trajectory gen}}\cdot\underbrace{\prod_{t}p(\mathbf{x}_{t}|\boldsymbol{\tau}_{t},\mathcal{W})}_{\text{sensor simulation}}.$
@@
Figure 3: Overview of the LidarDM: Given the input traffic layout at time $t=0$, LidarDM begins by generating actors and the intensity-infused static scene. We then generate the motion of the actors and the egocar, and compose the underlying 4D world. Finally, a generative- and physics-based simulation is used to create realistic 4D sensor data.
III-B Scene, Object and Trajectory Generation
We decompose the world into a static, intensity-infused background scene that is constant over time and dynamic foreground objects that move. This decomposition simplifies the challenging 4D world generation into manageable tasks: creating object geometries and generating dynamic effects, while ensuring temporal consistency (e.g. constant cars’ shapes and walls and trees over time) and physical plausibility (e.g. ensuring correct collision reasoning).
III-B1 Scene Generation
The scene generation step addresses the problem of sampling the geometry and LiDAR intensity of a scene from a given input layout $\mathcal{I}$: $\mathbf{s}\sim p(\mathbf{s}|\mathcal{I})$. We parameterize the intensity-infused scene $\mathbf{s}\in\mathbb{R}^{2\times L\times W\times H}$, where each entry $s_{j}\in\mathbb{R}^{2}$ encodes the truncated signed distance to the surface and an intensity value.
Model: We leverage the latent diffusion model [2, 2] to tackle this challenge of modeling and sampling from $p(\mathbf{s}|\mathcal{I})$ due to its capacity to sample high-quality data while effectively incorporating strong conditional guidance. Specifically, our model encodes the high-dimensional $\mathbf{s}$ into a continuous latent representation $\mathbf{z}$ using an encoder-decoder structure [2] with a scene encoder $E_{\theta}(\mathbf{s})=\mathbf{z}$ and a scene decoder $D_{\theta}(\mathbf{z})=\tilde{\mathbf{s}}$. This encoder-decoder structure efficiently compresses the input data into a lower-dimensional latent space, enabling more effective and efficient sampling. Additionally, we encode our high-definition map layout $\mathcal{I}$ into a latent space $\mathbf{c}=M_{\theta}(\mathcal{I})$, allowing for more compact conditioning.